# Asymptotic spreading of predator-prey populations in a shifting environment

### v1

@@ 1 Introduction
A fundamental challenge in ecology is to determine the factors leading to the persistence and spread of a given species in an environment. These issues are receiving a renewed interest as changes in climatic conditions can dramatically impact the suitability of a habitat for a species’ survival and growth. As temperatures rise, many species have moved in the directions of the poles or toward higher elevations, in an apparent attempt to keep pace with shifting temperature isotherms [54, 43]. As species establish in new regions, new biotic interactions take place [29], which in turn can have significant consequences for species abundance and biodiversity [62, 48, 60], the functioning of ecosystems [60, 61, 67], the spread of disease [55, 70] and human welfare [55]. The ecological effects of a changing climate are complex and various. While many species are vulnerable to a changing climate, for many others climate-related changes may facilitate expansion to new areas and population growth [63, 3, 10, 33].
Mathematical modeling can be used to determine why certain species decline while others prosper under the changing climate. The study of species persistence and spread often depend on the spatial context and much analysis in the classical literature has been based on reaction-diffusion models, of which a prominent example is the Fisher-KPP equation [26, 38], which describes the spreadin of a single population
@@
where $d$ corresponds to the dispersal rate and $r$ the intrinsic growth rate of a species $u$. For a given population density $u(t,x)$, Aronson and Weinberger introduced the key notion of spreading speed, which refers to the number $c^{*}>0$ such that
$\lim_{t\to\infty}\sup_{|x|>ct}u(t,x)=0\quad\text{ for }c\in(c^{*},\infty),\quad\text{ and }\quad\lim_{t\to\infty}\inf_{|x|>ct}u(t,x)>0\quad\text{ for }c\in(0,c^{*}).$ (2)
The problem of spreading speed for the Fisher-KPP equation was first investigated by Kolmogorov et al [38] for heaviside initial condition, who showed that
$c^{*}=2\sqrt{f^{\prime}(0)}.$
This was later generalized to any compactly supported initial data and in higher spatial dimensions by Aronson and Weinberger [2]. This theoretical spreading speed has yielded good estimates for range expansion observed in nature [58].
@@
A shifting environment also leads to new spreading phenomena. In [40], the spreading speed for solutions of Fisher’s equation with growth rate $f(u,x-ct)=(r(x-ct)-u)u$ was determined using the Hamilton-Jacobi method, in the case that the intrinsic growth rate of the species is positive and monotone. They showed that, for a certain range of velocities of climate shift, the species spreads with speed distinct from either of the limiting KPP invasion speeds in a phenomenon called non-local pulling [34, 30]. When the growth rate is non-monotone, the existence of forced wave and their attractivity is studied in [8].
@@ 1.1 The predator-prey model in a shifting environment
To this end, we propose the following predator-prey model of reaction-diffusion type to analyze the consequence of such effects on the population dynamics and spread of species.
@@
By observing (via maximum principle) that the density of the predator is bounded from above by $\|a\|_{\infty}-1$, it follows that the quantity $\beta\coloneqq 1-b(\left\lVert a\right\rVert_{\infty}-1)$ corresponds to the maximum carrying capacity for the prey.
@@ 1.2 Main Results
Following [32, Definition 1.2], we introduce the closely related notion of maximal and minimal speed $\overline{c}_{*}$, $\underline{c}_{*}$:
@@
Let $c_{1}>0$ be given, $a:\mathbb{R}\to\mathbb{R}$ be increasing, and suppose (H1) holds. If $(u(t,x),v(t,x))$ is the solution of (4) with initial data satisfying (IC), then spreading speed of $u$ exists, and is given by
@@
Let $c_{1}>0$ be given, $a:\mathbb{R}\to\mathbb{R}$ be decreasing, and suppose (H1) holds. Let $u(t,x)$ be the solution of (4). Then the spreading speed of $u$ exists, and is given by
$c^{*}=\begin{cases}2\sqrt{r_{2}}&\text{ if }c_{1}\leq 2\sqrt{r_{2}},\\ 2\sqrt{r_{1}}&\text{ if }c_{1}\geq 2\sqrt{r_{1}}.\end{cases}$
@@
We also show the convergence of $(u,v)$ to the homogeneous state in the moving frame with speed different from $c_{1}$ and $c^{*}$.
@@
    If $c\in(c^{*},\infty)$, then $(u_{c},v_{c})\to(0,1)$ in $C_{loc}({\mathbb{R}^{2}}).$
@@
    If $c\in(0,c^{*})\setminus\{c_{1}\}$, then $(u_{c},v_{c})\to(\tilde{u},\tilde{v})$ in $C_{loc}(\mathbb{R}^{2})$, where $(\tilde{u},\tilde{v})$ is the unique positive equilibrium of the kinetic system
@@
Note that the case $c_{1}\in(2\sqrt{r_{2}},2\sqrt{r_{1}})$ is not covered by Theorem 2. In that case the Hamilton-Jacobi approach does not apply. We conjecture that $c^{*}=c_{1}$ in that case and the predator advances in locked step with the environment. See [8, 25] for results regarding a single species in a shifting habitat. A possible approach is to use the persistence theory as in [20].
@@
In the case of $a(\cdot)\equiv a_{0}$ being a constant and $v_{0}\equiv 1$, the spreading speed of the predator was determined by Pan in [53, Theorem 2.1].
To consolidate the formulas for the spreading speeds in Theorems 1 and 2, we will denote $\lambda^{*}=\frac{c_{1}}{2}-\sqrt{|r_{2}-r_{1}|}$, $r_{1}=a(-\infty)-1$, $r_{2}=a(+\infty)-1$. Then the spreading speed for all cases can be given by
$\sigma(c_{1};r_{1},r_{2})=\begin{cases}2\sqrt{r_{2}}&\text{ if }r_{1}<r_{2}\quad\text{and}\quad c_{1}\leq 2\sqrt{r_{2}}\\ \lambda^{*}+\frac{r_{1}}{\lambda^{*}}&\text{ if }r_{1}<r_{2}\quad\text{and}\quad 2\sqrt{r_{2}}<c_{1}<2\sqrt{r_{1}}+2\sqrt{r_{2}-r_{1}}\\ 2\sqrt{r_{1}}&\text{ if }r_{1}<r_{2}\quad\text{and}\quad c_{1}\geq 2\sqrt{r_{1}}+2\sqrt{r_{2}-r_{1}},\\ 2\sqrt{r_{2}}&\text{if }r_{1}>r_{2}\quad\text{and}\quad c_{1}\leq 2\sqrt{r_{2}}\\ 2\sqrt{r_{1}}&\text{if }r_{1}>r_{2}\quad\text{and}\quad c_{1}\geq 2\sqrt{r_{1}},\end{cases}$ (9)
@@
$\sigma(c_{1};r_{1},r_{2})=\begin{cases}2\sqrt{r_{2}}&\text{ if }c_{1}\leq 2\sqrt{r_{2}}\\ \lambda^{*}+\frac{r_{1}}{\lambda^{*}}&\text{ if }r_{1}<r_{2}\quad\text{and}\quad 2\sqrt{r_{2}}<c_{1}<2\sqrt{r_{1}}+2\sqrt{r_{2}-r_{1}}\\ 2\sqrt{r_{1}}&\text{ if }c_{1}\geq 2\sqrt{r_{1}}+2\sqrt{\max\{0,r_{2}-r_{1}\}},\end{cases}$ (10)
@@ 1.3 Related mathematical results
We also mention a closely related work of Choi, Giletti, and Guo [14], where they considered a two-species predator-prey system similar to (4), with the intrinsic growth rate $r=r(x-c_{1}t)$ for the prey subject to the climate shift instead of the coefficient $a$. They considered the case when both initial data $u_{0}$ and $v_{0}$ are compactly supported and a non-decreasing profile for the growth rate with $r$ changing sign, $r(-\infty)<0<r(\infty)$. In the case of local dispersal, they showed that the prey persists by spreading if and only if the maximal speed of the prey exceeds the environmental speed (i.e., $2\sqrt{dr(+\infty)}>c_{1}$), while the predator persists by spreading at the speed given by the smaller of the prey and maximal predator spreading speeds. In their setting both species tends to zero in $\{(t,x):x<c_{1}t\}$, while in the zone ahead of the environmental shift, the density of the prey is strictly decreasing so there is no nonlocally pulling phenomenon. We also mention [31] for the case of two weak-competing predators and one prey, and [1], for the case of one predator and two preys. For compactly supported initial data, the invasion wave of the prey resembles the effect of a shifting environment studied in our paper. However, the exact spreading speed of the predator(s) is not completely determined.
@@ 2 Upper bound on spreading speed
In this section, we give a quick proof of $\overline{c}_{*}\leq\sigma(c_{1};r_{1},r_{2})$, where $\sigma(c_{1};r_{1},r_{2})$ is given by the first three cases of (9). i.e. the spreading speed $c^{*}$ is bounded above by $\sigma(c_{1};r_{1},r_{2})$.
@@
By the classical theory of reaction-diffusion equations, there exists a unique solution $(u,v)$ satisfying (4) for all $(t,x)\in(0,\infty)\times\mathbb{R}$; see, e.g. [59]. Moreover, since $10\leq u_{0}\leq\left\lVert a\right\rVert_{\infty}-$ 1 and $\beta\leq v_{0}\leq 1$, the maximum principle (see [57, Chapter 3, Section 6, Theorem 10] or [39, Theorem 6.2.1]) implies $10\leq u\leq\left\lVert a\right\rVert_{\infty}-$ 1 and $\beta\leq v\leq 1$ on $(0,\infty)\times\mathbb{R}$. ∎
@@
In the case $r_{2}>r_{1}$, we may invoke [40, Theorem 6] to deduce that $\bar{u}$ satisfies
@@
$U(t,x)=\begin{cases}\exp(-\sqrt{r_{1}}+2r_{1}t)&\text{ if }c_{1}\geq 2\sqrt{r_{1}},\\ \min\{1,\exp(-\lambda(x-c_{1}t))\}&\text{ if }2\sqrt{r_{2}}<c_{1}<2\sqrt{r_{1}},\leavevmode\nobreak\ \lambda=\tfrac{1}{2}\left(c_{1}-\sqrt{c_{1}^{2}-4r_{2}}\right)\\ \min\{1,\exp(-\sqrt{r_{2}}x+2r_{2}t\}&\text{ if }0<c_{1}\leq 2\sqrt{r_{2}}.\end{cases}$
Then it can be verified that $\overline{u}$ is a generalized supersolution of (12) (see [39, Definition 1.1.1] or [9, Definition 4.2] for the definition). Hence, we again deduce that (14) holds where
$\sigma=\begin{cases}2\sqrt{r_{2}}&\text{ if }c_{1}\leq 2\sqrt{r_{2}},\\ 2\sqrt{r_{1}}&\text{ if }c_{1}\geq 2\sqrt{r_{1}},\\ c_{1}&\text{ otherwise.}\end{cases}$
@@ 3 Rough Estimate for $v(t,x)$
Since $v(t,x)\leq 1$ (thanks to Lemma 1), it suffices to show the lower bound. We shall follow the proof of Theorem 5.1 in [20]. Fix $c>\sigma(c_{1};r_{1},r_{2})$. we may suppose for contradiction that there exists a sequence $\{(t_{n},x_{n})\}$ with $t_{n}\to\infty$ and $x_{n}\geq ct_{n}$ such that $\limsup\limits_{n\to\infty}v(t_{n},x_{n})<1$. Denote $(u_{n},v_{n})(t,x)=(u,v)(t+t_{n},x+x_{n})$. By standard parabolic estimates, we may pass to a further subsequence so that $(u_{n},v_{n})$ converges to an entire in time solution $(u_{\infty},v_{\infty})$ of (4) in $C_{loc}(\mathbb{R}^{2})$. Since $c>\sigma(c_{1};r_{1},r_{2})$, by Lemma 1, we have $u_{\infty}\equiv 0$. Thus, $v_{\infty}$ is an entire solution satisfying the equation
@@
Since $v\geq\beta$ for all $(t,x)\in(0,\infty)\times\mathbb{R}$, we deduce that $v_{\infty}\geq\beta$ for all $(t,x)\in\mathbb{R}^{2}$. By the classification of entire solution of the diffusive logistic equation (see, e.g. [49, Lemma 2.3(d)]) we have $v_{\infty}\equiv 1$. This is in contradiction with the statement $\limsup\limits_{n\to\infty}v(t_{n},x_{n})<1$. ∎
@@
$R_{1}(s)=\begin{cases}R_{1a}(s)=\begin{cases}r_{2}&\text{for }s\geq 2\sqrt{r_{2}}\\ \underline{r}_{2}&\text{for }c_{1}\leq s<2\sqrt{r_{2}}\\ \underline{r}_{1}&\text{for }s<c_{1}\end{cases}&\quad\text{if }c_{1}<2\sqrt{r_{2}},\\ R_{1b}(s)=\begin{cases}r_{2}&\text{for }s\geq c_{1},\\ r_{1}&\text{for }\lambda+\frac{r_{1}}{\lambda}\leq s<c_{1}\\ \underline{r}_{1}&\text{for }s<\lambda+\frac{r_{1}}{\lambda}\end{cases}&\quad\text{if }2\sqrt{r_{2}}<c_{1}<2(\sqrt{r_{1}}+\sqrt{r_{2}-r_{1}}),\\ R_{1c}(s)=\begin{cases}r_{2}&\text{for }s\geq c_{1},\\ r_{1}&\text{for }2\sqrt{r_{1}}\leq s<c_{1}\\ \underline{r}_{1}&\text{for }s<2\sqrt{r_{1}}\end{cases}&\quad\text{if }c_{1}>2(\sqrt{r_{1}}+\sqrt{r_{2}-r_{1}})\end{cases}$ (19)
In Cases 2(a)-(b), we have $r_{1}>r_{2}$, and we let
$R_{2}(s)=\begin{cases}R_{2a}(s)=\begin{cases}r_{2}&\text{for }s\geq 2\sqrt{r_{2}},\\ \underline{r}_{2}&\text{for }c_{1}\leq s<2\sqrt{r_{2}}\\ \underline{r}_{1}&\text{for }s<c_{1}.\end{cases}&\quad\text{if }c_{1}<2\sqrt{r_{2}}\\ R_{2b}(s)=\begin{cases}r_{2}&\text{for }s\geq c_{1},\\ r_{1}&\text{for }2\sqrt{r_{1}}\leq s<c_{1}\\ \underline{r}_{1}&\text{for }s<2\sqrt{r_{1}}\end{cases}&\quad\text{if }c_{1}>2\sqrt{r_{1}}.\end{cases}$ (20)

### v2

@@ 1 Introduction
A fundamental challenge in ecology is to understand the persistence and spread of a given species in an environment. These issues are receiving a renewed interest as changes in climatic conditions can dramatically impact the suitability of a habitat for a species’ survival and growth. As temperatures rise, many species have moved in the directions of the poles or toward higher elevations, in an apparent attempt to keep pace with shifting temperature isotherms [55, 44]. As species establish in new regions, new biotic interactions take place [30], which in turn can have significant consequences for species abundance and biodiversity [63, 49, 61], the functioning of ecosystems [61, 62, 68], the spread of disease [56, 71] and human welfare [56]. The ecological effects of a changing climate are complex and various. While many species are vulnerable to a changing climate, for many others climate-related changes may facilitate expansion to new areas and population growth [64, 4, 11, 34].
Mathematical modeling can be used to determine why certain species decline while others prosper under the changing climate. The study of species persistence and spread often depend on the spatial context, and much analysis in the classical literature has been based on reaction-diffusion models. A prominent example is the Fisher-KPP equation [27, 39], which describes the spreading of a single population
@@
where $d$ corresponds to the dispersal rate and $r$ the intrinsic growth rate of a species $u$. For a given population density $u(t,x)$, Aronson and Weinberger [2, 3] introduced the key notion of spreading speed, which refers to the number $c^{*}>0$ such that
$\lim_{t\to\infty}\sup_{|x|>ct}u(t,x)$ $=0\quad\text{ for }c\in(c^{*},\infty),\quad\text{and}\quad\lim_{t\to\infty}\inf_{|x|>ct}u(t,x)$ $>0\quad\text{ for }c\in(0,c^{*}).$
The problem of spreading speed for more general equations of the form
$u_{t}=du_{xx}+f(u)\quad\text{ for }x\in\mathbb{R},\leavevmode\nobreak\ t>0,$ (2)
was first investigated by Kolmogorov et al. [39] for heaviside initial condition, who showed under certain assumptions on the growth function $f(u)$ that
$c^{*}=2\sqrt{df^{\prime}(0)}.$
This result was later generalized to any compactly supported initial data and in higher spatial dimensions by Aronson and Weinberger [3]. This theoretical spreading speed has yielded good estimates for range expansion observed in nature [59].
@@
A shifting environment also leads to new spreading phenomena. In [41], the spreading speed for solutions of Fisher’s equation with growth rate $f(u,x-ct)=(r(x-ct)-u)u$ was determined using the Hamilton-Jacobi method, in the case that the intrinsic growth rate of the species is positive and monotone. They showed that, for a certain range of velocities of climate shift, the species spreads with speed distinct from either of the limiting KPP invasion speeds in a phenomenon called non-local pulling [35, 31]. When the growth rate is non-monotone, the existence of forced waves and their attractivity is studied in [9].
@@ 1.1 The predator-prey model in a shifting environment
To this end, we propose the following predator-prey model of reaction-diffusion type to analyze the consequence of such effects on the population dynamics and spread of species:
@@
By observing (via maximum principle) that the density of the predator is bounded from above by $\|a\|_{\infty}-1$, it follows that the quantity $\beta\coloneqq 1-b(\left\lVert a\right\rVert_{\infty}-1)$ corresponds to the minimum carrying capacity for the prey.
@@ 1.2 Main Results
Following [33, Definition 1.2], we introduce the closely related notion of maximal and minimal speed $\overline{c}_{*}$, $\underline{c}_{*}$:
@@
Let $c_{1}>0$ be given, $a:\mathbb{R}\to\mathbb{R}$ be increasing, and suppose (H1) holds. If $(u(t,x),v(t,x))$ is the solution of (4) with initial data satisfying (IC), then the spreading speed of $u$ exists, and is given by
@@
Let $c_{1}\in\mathbb{R}\setminus(2\sqrt{r_{2}},2\sqrt{r_{1}}]$ be given, $a:\mathbb{R}\to\mathbb{R}$ be decreasing, and suppose (H1) holds. If $(u(t,x),v(t,x))$ is the solution of (4) with initial data satisfying (IC), then the spreading speed of $u$ exists, and is given by
$c^{*}=\begin{cases}2\sqrt{r_{2}}&\text{ if }c_{1}\leq 2\sqrt{r_{2}},\\ 2\sqrt{r_{1}}&\text{ if }c_{1}>2\sqrt{r_{1}}.\end{cases}$
@@
We also show the convergence of $(u,v)$ to the homogeneous state in the moving frames with speed different from $c_{1}$ or $c^{*}$.
@@
    If $c\in(c^{*},\infty)$, then $\lim_{t\to\infty}(u_{c}(\cdot,t),v_{c}(\cdot,t))\to(0,1)$ in $C_{loc}({\mathbb{R}})$.
@@
    If $c\in(0,c^{*})\setminus\{c_{1}\}$, then
    $\lim_{t\to\infty}(u_{c}(\cdot,t),v_{c}(\cdot,t))\to(\tilde{u},\tilde{v})\quad\text{ in }C_{loc}({\mathbb{R}}),$
    where $(\tilde{u},\tilde{v})$ is the unique positive equilibrium of the kinetic system
@@
Note that the case $c_{1}\in(2\sqrt{r_{2}},2\sqrt{r_{1}}]$ is not covered by Theorem 2. In that case the Hamilton-Jacobi approach does not directly apply. We conjecture that $c^{*}=c_{1}$ in that case and the predator advances in locked step with the environment. See [9, 26] for results regarding a single species in a shifting habitat. A possible approach is to use the persistence theory as in [21].
@@
In the case of $a(\cdot)\equiv a_{0}$ being a constant and $v_{0}\equiv 1$, the spreading speed of the predator was determined by Pan in [54, Theorem 2.1].
To consolidate the formulas for the spreading speeds in Theorems 1 and 2, we will denote $\lambda^{*}=\frac{c_{1}}{2}-\sqrt{|r_{2}-r_{1}|}$. Then the spreading speed for all cases can be given by
$\sigma(c_{1};r_{1},r_{2})=\begin{cases}2\sqrt{r_{2}}&\text{ if }r_{1}<r_{2}\quad\text{and}\quad c_{1}\leq 2\sqrt{r_{2}}\\ \lambda^{*}+\frac{r_{1}}{\lambda^{*}}&\text{ if }r_{1}<r_{2}\quad\text{and}\quad 2\sqrt{r_{2}}<c_{1}<2\sqrt{r_{1}}+2\sqrt{r_{2}-r_{1}}\\ 2\sqrt{r_{1}}&\text{ if }r_{1}<r_{2}\quad\text{and}\quad c_{1}\geq 2\sqrt{r_{1}}+2\sqrt{r_{2}-r_{1}},\\ 2\sqrt{r_{2}}&\text{ if }r_{1}>r_{2}\quad\text{and}\quad c_{1}\leq 2\sqrt{r_{2}}\\ 2\sqrt{r_{1}}&\text{ if }r_{1}>r_{2}\quad\text{and}\quad c_{1}>2\sqrt{r_{1}},\end{cases}$ (10)
@@
$\sigma(c_{1};r_{1},r_{2})=\begin{cases}2\sqrt{r_{2}}&\text{ if }c_{1}\leq 2\sqrt{r_{2}}\\ \lambda^{*}+\frac{r_{1}}{\lambda^{*}}&\text{ if }r_{1}<r_{2}\quad\text{and}\quad 2\sqrt{r_{2}}<c_{1}\leq 2\sqrt{r_{1}}+2\sqrt{r_{2}-r_{1}}\\ 2\sqrt{r_{1}}&\text{ if }c_{1}>2\sqrt{r_{1}}+2\sqrt{\max\{0,r_{2}-r_{1}\}},\end{cases}$ (11)
@@ 1.3 Related mathematical results
We also mention a closely related work of Choi, Giletti, and Guo [15], where they considered a two-species predator-prey system similar to (4), with the intrinsic growth rate $r=r(x-c_{1}t)$ for the prey subject to the climate shift instead of the coefficient $a$. They considered the case when both initial data $u_{0}$ and $v_{0}$ are compactly supported and a non-decreasing profile for the growth rate with $r$ changing sign, $r(-\infty)<0<r(\infty)$. In the case of local dispersal, they showed that the prey persists by spreading if and only if the maximal speed of the prey exceeds the environmental speed (i.e., $2\sqrt{dr(+\infty)}>c_{1}$), while the predator persists by spreading at the speed given by the smaller of the prey and maximal predator spreading speeds. In their setting both species tends to zero in $\{(t,x):x<c_{1}t\}$, while in the zone ahead of the environmental shift, the density of the prey is strictly decreasing so there is no nonlocal pulling phenomenon. We also mention [32] for the case of two weak-competing predators and one prey, and [1], for the case of one predator and two preys. For compactly supported initial data, the invasion wave of the prey resembles the effect of a shifting environment studied in our paper. However, the exact spreading speed of the predator(s) is not completely determined.
@@ 2 Upper bound on spreading speed
In this section, we give a quick proof of $\overline{c}_{*}\leq\sigma(c_{1};r_{1},r_{2})$, where $\sigma(c_{1};r_{1},r_{2})$ is given by the first three cases of (10), i.e., the spreading speed $c^{*}$ is bounded above by $\sigma(c_{1};r_{1},r_{2})$.
@@
By the classical theory of reaction-diffusion equations, there exists a unique solution $(u,v)$ satisfying (4) for all $(t,x)\in(0,\infty)\times\mathbb{R}$; see, e.g. [60]. Moreover, since $10\leq u_{0}\leq\left\lVert a\right\rVert_{\infty}-$ 1 and $\beta\leq v_{0}\leq 1$, the maximum principle (see [58, Chapter 3, Section 6, Theorem 10] or [40, Theorem 6.2.1]) implies $10\leq u\leq\left\lVert a\right\rVert_{\infty}-$ 1 and $\beta\leq v\leq 1$ on $(0,\infty)\times\mathbb{R}$. ∎
@@
In the case $r_{2}>r_{1}$, we may invoke [41, Theorem 6] to deduce that $\bar{u}$ satisfies
@@
$U(t,x)=\begin{cases}\exp(-\sqrt{r_{1}}{x}+2r_{1}t)&\text{ if }c_{1}\geq 2\sqrt{r_{1}},\\ \min\{r_{1},\exp(-\lambda(x-c_{1}t))\}&\text{ if }2\sqrt{r_{2}}<c_{1}<2\sqrt{r_{1}},\leavevmode\nobreak\ \lambda=\tfrac{1}{2}\left(c_{1}-\sqrt{c_{1}^{2}-4r_{2}}\right)\\ \min\{r_{1},\exp(-\sqrt{r_{2}}x+2r_{2}t)\}&\text{ if }0<c_{1}\leq 2\sqrt{r_{2}}.\end{cases}$
Then it can be verified that $U$ is a generalized supersolution of (13) (see [40, Definition 1.1.1] or [10, Definition 4.2] for the definition). Hence, we again deduce that (15) holds where
$\sigma=\begin{cases}2\sqrt{r_{1}}&\text{ if }c_{1}\geq 2\sqrt{r_{1}},\\ 2\sqrt{r_{2}}&\text{ if }c_{1}\leq 2\sqrt{r_{2}},\\ c_{1}&\text{ otherwise.}\end{cases}$
@@ 3 Rough Estimate for $v(t,x)$
Since $v(t,x)\leq 1$ (thanks to Lemma 1), it suffices to show the lower bound. We shall follow the proof of Theorem 5.1 in [21]. Fix $c>\sigma(c_{1};r_{1},r_{2})$. we may suppose for contradiction that there exists a sequence $\{(t_{n},x_{n})\}$ with $t_{n}\to\infty$ and $x_{n}\geq ct_{n}$ such that $\limsup\limits_{n\to\infty}v(t_{n},x_{n})<1$. Denote $(u_{n},v_{n})(t,x)=(u,v)(t+t_{n},x+x_{n})$. By standard parabolic estimates, we may pass to a further subsequence so that $(u_{n},v_{n})$ converges to an entire in time solution $(u_{\infty},v_{\infty})$ of (4) in $C_{loc}(\mathbb{R}^{2})$. Since $c>\sigma(c_{1};r_{1},r_{2})$, by Proposition 1, we have $u_{\infty}\equiv 0$. Thus, $v_{\infty}$ is an entire solution satisfying the equation
@@
Since $v\geq\beta$ for all $(t,x)\in(0,\infty)\times\mathbb{R}$, we deduce that $v_{\infty}\geq\beta$ for all $(t,x)\in\mathbb{R}^{2}$. By the classification of entire solution of the diffusive logistic equation (see, e.g. [50, Lemma 2.3(d)]) we have $v_{\infty}\equiv 1$. This is in contradiction with the statement $\limsup\limits_{n\to\infty}v(t_{n},x_{n})<1$. ∎
@@
$R_{1}(s)=\begin{cases}R_{1a}(s)=\begin{cases}r_{2}&\text{for }s>2\sqrt{r_{2}}\\ \underline{r}_{2}&\text{for }c_{1}<s\leq 2\sqrt{r_{2}}\\ \underline{r}_{1}&\text{for }s\leq c_{1}\end{cases}&\quad\text{if }c_{1}\leq 2\sqrt{r_{2}},\\ R_{1b}(s)=\begin{cases}r_{2}&\text{for }s>c_{1},\\ r_{1}&\text{for }\lambda^{*}+\frac{r_{1}}{\lambda^{*}}<s\leq c_{1}\\ \underline{r}_{1}&\text{for }s\leq\lambda^{*}+\frac{r_{1}}{\lambda^{*}}\end{cases}&\quad\text{if }2\sqrt{r_{2}}<c_{1}<2(\sqrt{r_{1}}+\sqrt{r_{2}-r_{1}}),\\ R_{1c}(s)=\begin{cases}r_{2}&\text{for }s>c_{1},\\ r_{1}&\text{for }2\sqrt{r_{1}}<s\leq c_{1}\\ \underline{r}_{1}&\text{for }s\leq 2\sqrt{r_{1}}\end{cases}&\quad\text{if }c_{1}\geq 2(\sqrt{r_{1}}+\sqrt{r_{2}-r_{1}})\end{cases}$ (19)
In Cases 2(a)-(c), we have $r_{1}>r_{2}$, and we let
$R_{2}(s)=\begin{cases}R_{2a}(s)=\begin{cases}r_{2}&\text{for }s>2\sqrt{r_{2}},\\ \underline{r}_{2}&\text{for }c_{1}\leq s\leq 2\sqrt{r_{2}}\\ \underline{r}_{1}&\text{for }s<c_{1}.\end{cases}&\quad\text{if }c_{1}<2\sqrt{r_{2}},\\ R_{2b}(s)=\begin{cases}r_{2}&\text{for }s>2\sqrt{r_{2}},\\ \min\{r_{2},\underline{r}_{1}\}&\text{for }s=2\sqrt{r_{2}}\\ \underline{r}_{1}&\text{for }s<2\sqrt{r_{2}}.\end{cases}&\quad\text{if }c_{1}=2\sqrt{r_{2}},\\ R_{2c}(s)=\begin{cases}r_{2}&\text{for }s\geq c_{1},\\ r_{1}&\text{for }2\sqrt{r_{1}}<s<c_{1}\\ \underline{r}_{1}&\text{for }s\leq 2\sqrt{r_{1}}\end{cases}&\quad\text{if }c_{1}>2\sqrt{r_{1}}.\end{cases}$ (20)